# Field-free, Quasi-continuous Operation of Optical Nanofiber Interface with Two-dimensional Ferromagnetic Trap

### v3

@@ Abstract
A soft ferromagnetic foil uniformizes Tesla-level magnetic fields generated by attached permanent magnets, producing a uniform and electronically tunable surface field on the opposite side. By arranging $n$ precisely fabricated rectangular foils, a nearly ideal magnetic quadrupole field with a substantial gradient can be created at center. This robust and tunable field configuration is useful for 2-dimensional magneto-optical trapping (2D-MOT) and magnetic guiding of cold atoms. In this work, by aligning an optical nanofiber (ONF) to the zero-field line of a 2-foil-based planar 2D-MOT, we demonstrate field-free operation of the quantum optical interface in a quasi-continuous manner, without switching off the magnetic field. Transient transmission spectroscopy is performed with a measurement repetition rate as high as 250 kHz. An anomalous line broadening is observed, which is not fully understood, but is partly explained by a small residual field along the zero-field line. Through additional field measurements and simulations, we clarify that this residual field can be eliminated in an $n$=4 assembly, resulting in an ultra-straight 2D trap to support efficient sub-Doppler cooling and uniform light-atom interaction over exceptionally long field-free distances $l$. With the strong field gradient to support atom guiding, the ferromagnetic device may also enable new quantum optical scenarios featuring interactions between co-guided atoms and photons.
@@ I Introduction
A potential solution to meet the field-free requirements without switching off the field is to load the ONF interface with a two-dimensional MOT [18]. As in Fig. 1a, 2D-MOT operates with magneto-optical trapping force in the $x-y$ plane only, without requiring any field gradient along $z$. Cold atoms trapped in 2D-MOT is highly elongated, naturally serve to load the ONF interface. By precisely locating the ONF to the zero-field line, the magnetic field sensed by cold atoms in the nanoscale near field is practically zeroed, even for 2D-MOT operates with very high field gradients [19]. Importantly, since polarization gradient cooling is most efficient at zero field [20, 18], we expect efficient loading of atoms to the ONF optical lattice [10, 12, 13, 14]. However, all these 2D-MOT benefits require precise alignment of the zero-field line with the long ONF. While ONF is itself highly stiff [5] to support a straight line within e.g. a centimeter, a 2D-MOT optimized for free-space quantum optics [21] does not guarantee a straight enough zero-field line to match the straight ONF. In addition, to construct a 2D-MOT usually requires closely spaced rectangular coils carrying high currents [21], a setup that may be challenging to implement to ONF-based quantum optical platforms [22, 23, 24] where the ONF mounting structures necessarily constrains the size and complexity of the setup.
@@
Figure 1: (a): Schematic of an ONF-2D-MOT interface with interaction length $l$. (b): Uniformization of magnetic field by a rectangular ${\bf\mu}-$metal foil. For the precisely machined foil, the magnetic field within the dashed line box on top can be highly uniform, being nearly invariant along $z$. The center of the $S_{2}$ surface ${\bf r}_{s}$ is marked with a red triangle marker. (c): The prism 2D-MOT using the $n$=2 foil structure in this work. Local laser beam polarizations defined along $+x,+y$ axes are marked. (d): 2D-MOT using $n$=4 foil structure. Local laser beam helicity are defined along $+x,+y$ axes in (a,c), and along a pair of $45^{\circ}$ axes in (d). In (c,d) the blue disk in the MOT center represents ONF. Its sub-micron diameter is exaggerated for clarity.
@@
In this work, we draw inspiration from earlier developments in magnetically guided atom interferometry [19] and use ferromagnetic trap for field-free operation of ONF interfaces. As in Fig. 1(b), soft-ferromagnetic ${\bf\mu}-$metal foil poled by permanent magnets from one side can generate highly uniform surface field on the other [25, 26]. By properly arranging $n=2$ (Fig. 1c) or $n=4$ (Fig. 1d) foils, a zero-field line is formed near the symmetry center to support the 2D-MOT operation. This zero-field line is ultra-straight [19]. Its position can be finely adjusted via the electronic current surrounding the foils. By shifting this zero-field line to overlap with the straight ONF, the magnetic field is nullified microscopically to support field-free operation without switching off the magnetic field for the 2D-MOT. We experimentally demonstrate the key ingredients of our proposal with a compact two-foil ONF interface, featuring microsecond-level cooling-measurement cycle time. An anomalous line-broadening [10] in transient ONF spectroscopy is observed, which is not fully understood, but is partly attributed to a residual field gradient along the zero-field line direction $z$ (Fig. 2c) in the $n=2$ assembly. This $z$-gradient can be completely eliminated in the $n=4$ configuration (Fig. 2(a,b)), with which we discuss on the prospects for quasi-continuous, field-free operation of ONF-cold-atom interface with $l\geq$100 mm interaction distances, and to enable novel quantum optical research featuring strongly coupled photons with matterwave in their respective waveguides.
@@
In the following our work is presented in three sections. First, in Sec. II, we outline the general principles of constructing an ultra-straight ferromagnetic trap for 2D cooling and trapping and for integrating the device with an optical nanofiber. In Sec. III, we present our experimental demonstration on field-free operation of the ONF interface, in particular, on a quasi-continuous transient absorption spectroscopy measurement with a repetition rate as high as 250 kHz. In Sec. IV we clarify practical aspects of the ultra-straight 2D field generation and propose an ONF interface with up to $l=10~{}$cm field-free interaction length. We summarize this work in Sec. V with an outlook into prospects of combining the ONF-interface with the ferromagnetically generated, nearly ideal 2D trapping potential.
@@ II.1 Robust uniformization of magnetic field
As in Fig. 1b, we consider rectangular ${\bf\mu}-$metal foil with width $W$, length $L$, and thickness $d\ll W,L$. The foil is magnetized primarily by permanent magnets, e.g. a Neodymium Iron Boron (NdFeB) block, from the $S_{1}$ side. To provide electronic tunability, the foil is in addition poled by current-carrying coils surrounding the foil. The NdFeB magnet with size $a\approx d$ is small. In absence of the ${\bf\mu}-$metal, its field at distance $W$ away is effectively a dipole field which varies substantially over any $O(W)$ distance. However, as highlighted by the dash-lined box in Fig. 1b, the field flux after being guided through the ${\bf\mu}-$metal becomes uniform near the distant surface $S_{2}$. Here, ${\bf r}_{h}={\bf r}_{s}+(x_{h},y_{h},z_{h})$ with $|x_{h}|,|y_{h}|=O(h)\ll W$ from the foil surface, and $|z_{h}|<L/4$ to avoid edge effects.
@@
Here, we provide an intuitive explanation on the robust field uniformity, which is analogous to field uniformization by an isolated conductor in electrostatics [25]. Specifically, the ${\bf\mu}-$metal foil responses to the source field with a redistribution of effective magnetic charge. Similar to electrostatic shielding by a grounded conductor, the magnetic charge redistribution shields the inhomogeneous source field near $S_{1}$ from affecting the proximity of $S_{2}$. The shielding is therefore highly efficient for foil with large enough $W$, $L$, and $d$. Certainly, to keep the magnetic charge neutrality, there are extra charges in the ${\bf\mu}-$metal, which tend to distribute uniformly on the surface to minimize the field energy [25]. This surface charge picture becomes accurate near $S_{2}$ where $B\ll B_{\text{sat}}$ so that the ${\bf\mu}-$metal response is linear. In addition, with ${\bf\mu}_{r}\gg 1$ in the ${\bf\mu}-$metal, the ${\bf B}$ field in air (or vacuum) is perpendicular to the surface, according to the tangential ${\bf H}=\frac{1}{{\bf\mu}}{\bf B}$ continuity. As in electrostatics, the uniform surface charge leads to highly uniform ${\bf B}$ near the flat $S_{2}$ surface, insensitive to small variations of source fields or the ${\bf\mu}_{r}$ distribution. The uniformization is also effective against imperfect current-coil winding [19]. This effect is most pronounced for thin-wire winding at a sufficient distance from the $S_{2}$ surface, as illustrated in Fig. 2(c,d). While the shielding of source magnetic irregularities tends to degrade at large working distances, the shielding can be reinforced, if necessary, by constructing additional, ${\bf\mu}-$metal enclosures for the sources (Fig. 2d).
@@
Finally, the field distribution is sensitive to the surface shape of the ${\bf\mu}-$metal itself, which can be machined with micrometer precision, over large scale. Mesoscopically, surface corrugations with a characteristic length scale $\xi$ of micrometers are typically present in these foils, but the associated fringe fields decay exponentially along $y$ [27, 28]. Therefore, to mitigate the effects of fringe field irregularities, 2D trapping should be positioned at a distance $y_{h}=h\gg\xi$.
@@
Figure 2: Numerical simulation of the 2D-quadruple field [29, 30] for the 2-foil assembly, (a,b), according to Sec. III.1; and 4-foil assembly, (c,d), according to Sec. IV.1. See Appendix A for simulation details. The foil dimensions are not to scale. A.S.: Optional, additional shielding of source magnet variations for operating 2D-trap at large $h,h^{\prime}$ distances [31].
II.2 $n=2$ foil trap
@@
As in Fig. 1c, Fig. 2(a,b) and Fig. 3, we consider arranging $n=2$ foils symmetrically in the $x-z$ plane with a gap distance of $\Delta y=2h$. With the opposite magnetization, the fields generated by the two foils cancel at ${\bf r}_{0}=(0,0,0)$ to form a magnetic zero. Near the center, the field ${\bf B}=B_{1}^{x}x{\bf e}_{x}+B_{1}^{y}y{\bf e}_{y}+B_{1}^{z}z{\bf e}_{z}$, with gradients $B_{1}^{x}:B_{1}^{y}:B_{1}^{z}=(1-\varepsilon):-1:\varepsilon$. Here $\varepsilon\sim(h/L)^{2}$ when $\varepsilon\ll 1$. The field zero is therefore a three-dimensional (3D) one. The residual $B_{1}^{z}$ gradient is associated with the fact that the field contributions from the two foils add up along $z$ for off-center locations. Nevertheless, with an operation distance $\varepsilon=h/L\ll 1$, the field distribution becomes highly elongated along $z$ to mimic a 2D field, as suggested by the dashed-line box in Fig. 2(c,d), with an approximate zero-field line near the center.
II.3 $n=4$ foil trap
@@
As in Fig. 1d and Fig. 2(c,d), we now consider arranging $n=4$ foils symmetrically in the $x-y$ plane, with a gap distance of $\Delta y=2h$ and $\Delta x=2h^{\prime}$ (Fig. 2b). As in Fig. 2c, in contrast to the $n=2$ case, in the $n=4$ assembly the field variation along $z$ caused by each foil due to the finite $L$ are completely canceled at $x=y=0$. This results in a quadrupole field distribution, ${\bf B}=B_{1}y{\bf e}_{x}+B_{1}x{\bf e}_{y}$ at the center. Furthermore, with $h,h^{\prime}\ll W$, the gradient $B_{1}$ varies extremely slowly within e.g. $|z|<L/4$, as suggested by the dashed-line box in Fig. 2(c,d).
II.4 Electronic tuning
An important advantage of soft ferromagnets is that their magnetization is electronically tunable. As suggested by Fig. 1 and demonstrated with the Fig. 3f data, a slight change of currents surrounding the foils offsets the ${\bf\mu}-$metal magnetization. By adjusting the currents surrounding the selected foils in the $n-$foil assembly, the zero-field line can be shifted in a precise and repeatable manner. As an extension of previous ferromagnetic trapping techniques [18, 19], hybridizing the permanent magnetization with electronic tuning results in strong and adjustable quadruple field gradient $B_{1}$ at large distance $h$, without requiring high electronic currents. Of course, the disadvantage of our approach here is that one cannot switch off the permanent magnets. Fortunately, since the ONF-2D-MOT interface only samples the nanoscale interaction zone, one only needs to align the zero-field line to the ONF location in the $x-y$ plane, for the field-free operation of atomic physics and/or quantum optics experiments. The magnetic field does not need to be switched off globally.
@@ II.4 Electronic tuning
Figure 3: (a): Schematic of the ONF-2D-MOT apparatus in this work. (b): The ONF-2D-MOT interface through the vacuum window. The ONF is visible through Rayleigh scattering of a 200 $\rm{\bf\mu}W$ heating light at 843 nm. (c): The 2D-MOT (marked with red arrow) and its three mirror images. (d): Schematic of the ONF-2D-MOT in planar geometry aided by two pairs of right-angle prisms. (e): Time-dependent fluorescence count through ONF and effective atom number $N_{\rm eff}$ during the MOT loading. The fit model for loading the density-limited MOT [20] (Appendix B), with $\rho_{0}=4\times 10^{10}/{\rm cm}^{3}$ final density near ONF, suggests a MOT lifetime of $\kappa^{-1}=0.4~{}{\rm sec}$. (f): Steady-state fluorescence count as a function of 2D-MOT position, adjusted by the $x$-coil and the foil-coil currents.
@@
For the two-foil assembly, the easiest way to form the quasi-2D-MOT is to use a right-angle prism pair that retro-reflect a single beam (Fig. 1c). The confinement along the $z-$direction is provided by two additional beams, also with two prisms in this work (Fig. 3d). This planer arrangement can be conveniently integrated with an ONF mount [13] (Fig. 3(a-c)), as suggested by Fig. 1c. By adjusting the foil-coil current to shift the field along $y$, and an $x-$coil current to shift the field along $x$, the zero-field line is precisely adjusted in the $n=2$ assembly to overlap with ONF. Clearly, the two-foil configuration only supports a quasi-2D-MOT. Depending on the $\varepsilon=h/L$ ratio, the weak gradient along the ONF direction may limit the field-free ONF-interaction length $l$.
On the other hand, as illustrated with Fig. 2(c,d), a true 2D quadruple field can be formed with the $n=4$ foil assembly. Similar to Fig. 1c, one prefers a planer geometry for the 2D-MOT construction, which can then be integrated with a standard ONF-mount [13]. Here, with the field lines rotated by $45^{\circ}$ comparing to the $n=2$ case, a 2D mirror MOT [19] is likely preferred for optimal optical accesses. The ONF can be integrated on top of the mirror along $z$. The zero-field line can be shifted freely in 2D by adjusting, e.g., the currents surrounding the foil pair on the left of Fig. 2d.

### v5

# Ferromagnetic Traps for Quasi-Continuous Operation of Optical Nanofiber Interfaces

@@ Abstract
A soft ferromagnetic plate uniformizes Tesla-level fields generated by attached permanent magnets, producing a smooth and electronically tunable surface field on the opposite side. By arranging $n$ precisely fabricated rectangular plates, a nearly ideal magnetic quadrupole field with a substantial gradient can be created at center. This robust and rapidly tunable field configuration is well suited for two-dimensional magneto-optical trapping (2D-MOT) and magnetic guiding of cold atoms. By aligning an optical nanofiber (ONF) along the zero-field line of a planar 2D-MOT in a 2-plate assembly, we demonstrate quasi-continuous, field-free operation of the quantum optical interface without switching off the magnetic field. Transient transmission spectroscopy with nanosecond laser pulses is performed on the 87Rb D2 line at a measurement repetition rate as high as 250 kHz. The observed line broadening, while not yet fully understood, is partially attributed to residual magnetic fields in the $n=2$ assembly. Through additional measurements and simulations, we verify that these residual fields can be fully eliminated in an $n=4$ assembly, resulting in an ultra-straight 2D trap that supports uniform light–atom interaction over exceptionally long, field-free distances. We extend our discussion to $n=6$, $n=8$ designs with similar uniformity but multiple zero-field lines. With its strong gradient for magnetic trapping, the ferromagnetic devices also enable new quantum optical scenarios featuring interactions between co-guided atoms and photons at ONF interfaces.
@@ I Introduction
A potential solution to meet the field-free requirements without switching off the field is to load the ONF interface with a two-dimensional MOT [18]. As in Fig. 1a, 2D-MOT operates with magneto-optical trapping force in the xyx-y plane only, without requiring any field gradient along $z$. Cold atoms trapped in 2D-MOT is highly elongated, naturally serve to load the ONF interface. By precisely locating the ONF to the zero-field line, the magnetic field sensed by cold atoms in the nanoscale near field is practically zeroed, even for 2D-MOT operates with very high field gradients [19]. Importantly, since polarization gradient cooling is most efficient at zero field [20, 18], we expect efficient loading of atoms to the ONF optical lattice [10, 12, 13, 14]. However, all these 2D-MOT benefits require precise alignment of the zero-field line with the long ONF. While ONF is itself highly stiff [5] to support a straight line within e.g. a centimeter, a 2D-MOT optimized for free-space quantum optics [21] does not guarantee a straight enough zero-field line to match the straight ONF. In addition, to construct a 2D-MOT usually requires closely spaced rectangular coils carrying high currents [21], a setup that may be challenging to implement to ONF-based quantum optical platforms [22, 23, 24] where the ONF mounting structures necessarily constrains the size and complexity of the setup.
@@
Figure 1: (a): Schematic of an ONF-2D-MOT interface with interaction length $l$. (b): Uniformization of magnetic field by a rectangular $\bf\mu$}-metal plate. For the precisely machined plate, the magnetic field within the dashed line box on top can be highly uniform, being nearly invariant along $z$. The center of the $S_{2}$ surface ${\bf r}_{s}$ is marked with a red triangle marker. (c): The prism 2D-MOT using the nn=2 plate structure in this work. Local laser beam polarizations defined along +x,+y+x,+y axes are marked. (d): 2D-MOT using nn=4 plate structure. Local laser beam helicity are defined along +x,+y+x,+y axes in (a,c), and along a pair of $45^{\circ}$ axes in (d). In (c,d) the blue disk in the MOT center represents ONF. Its sub-micron diameter is exaggerated for clarity.
@@
In this work, we draw inspiration from earlier developments in magnetically guided atom interferometry [19] and use ferromagnetic trap for field-free operation of ONF interfaces. As in Fig. 1(b), soft-ferromagnetic $\bf\mu$}-metal plate poled by permanent magnets from one side can generate highly uniform surface field on the other [25, 26]. By properly arranging $n=2$ (Fig. 1c) or $n=4$ (Fig. 1d) plates, a zero-field line is formed near the symmetry center to support the 2D-MOT operation. This zero-field line is ultra-straight [19]. Its position can be finely adjusted via the electronic current surrounding the plates. By shifting this zero-field line to overlap with the straight ONF, the magnetic field is nullified microscopically to support field-free operation without switching off the magnetic field for the 2D-MOT. We experimentally demonstrate the key ingredients of our proposal with a compact two-plate ONF interface, featuring microsecond-level cooling-measurement cycle time. An anomalous line-broadening in transient ONF spectroscopy is observed, which is not fully understood (Appendix C), but is partly attributed to a residual field gradient along the zero-field line direction $z$ (Fig. 2a) in the $n=2$ assembly. This $z$-gradient can be completely eliminated in the $n=4$ configuration (Fig. 2(c,d)), with which we discuss on the prospects for quasi-continuous, field-free operation of ONF-cold-atom interface with $l\geq$100 mm interaction distances, and to enable novel quantum optical research featuring strongly coupled photons with matterwave in their respective waveguides.
@@
In the following our work is presented in three sections. First, in Sec. II, we outline the general principles of constructing an ultra-straight ferromagnetic trap for 2D cooling and trapping and for integrating the device with an optical nanofiber. In Sec. III, we present our experimental demonstration on field-free operation of the ONF interface, in particular, on a quasi-continuous transient absorption spectroscopy measurement with a repetition rate as high as 250 kHz. In Sec. IV we clarify practical aspects of the ultra-straight 2D field generation and propose an ONF interface with up to l=10l=10~cm field-free interaction length. We summarize this work in Sec. V with an outlook into prospects of combining the ONF-technology with the ferromagnetically generated, nearly ideal 2D trapping potentials, particularly for interfacing co-guided photons and atoms.
@@ II.1 Robust uniformization of magnetic field
As in Fig. 1b, we consider rectangular $\bf\mu$}-metal plate with width $W$, length $L$, and thickness $d\ll W,L$. The plate is magnetized primarily by permanent magnets, e.g. a Neodymium Iron Boron (NdFeB) block, from the $S_{1}$ side. To provide electronic tunability, the plate is in addition poled by current-carrying coils surrounding the plate (see e.g. Fig. 2a). The NdFeB magnet with size $a\approx d$ is small. In absence of the $\bf\mu$}-metal, its field at distance $W$ away is effectively a dipole field which varies substantially over any $O(W)$ distance. However, as highlighted by the dash-lined box in Fig. 1b, the field flux after being guided through the $\bf\mu$}-metal becomes uniform near the distant surface $S_{2}$. Here, ${\bf r}_{h}={\bf r}_{s}+(x_{h},y_{h},z_{h})$ with $|x_{h}|,|y_{h}|=O(h)\ll W$ from the plate surface, and $|z_{h}|<L/4$ to avoid edge effects.
@@
Here, we provide an intuitive explanation on the robust field uniformity, which is analogous to field uniformization by an isolated conductor in electrostatics [25]. Specifically, the $\bf\mu$}-metal plate responses to the source field with a redistribution of effective magnetic charge. Similar to electrostatic shielding by a grounded conductor, the magnetic charge redistribution shields the inhomogeneous source field near $S_{1}$ from affecting the proximity of $S_{2}$. The shielding is therefore highly efficient for plate with large enough $W$, $L$, and $d$. Certainly, to keep the magnetic charge neutrality, there are extra charges in the $\bf\mu$}-metal, which tend to distribute uniformly on the surface to minimize the field energy [25]. This surface charge picture becomes accurate near $S_{2}$ where $B\ll B_{\text{sat}}$ so that the $\bf\mu$}-metal response is linear. In addition, with ${\bf\mu}_{r}\gg 1$ in the $\bf\mu$}-metal, the $\bf B$} field in air (or vacuum) is perpendicular to the surface, according to the tangential ${\bf H}=\frac{1}{{\bf\mu}}{\bf B}$ continuity. As in electrostatics, the uniform surface charge leads to highly uniform $\bf B$} near the flat $S_{2}$ surface, insensitive to small variations of source fields or the ${\bf\mu}_{r}$ distribution. The uniformization is also effective against imperfect current-coil winding [19]. This effect is most pronounced for thin-wire winding at a sufficient distance from the $S_{2}$ surface, as illustrated in Fig. 2(c,d). While the shielding of source magnetic irregularities tends to degrade at large working distances, the shielding can be reinforced, if necessary, by constructing additional, $\bf\mu$}-metal enclosures for the sources (Fig. 2d).
@@
Finally, the field distribution is sensitive to the surface shape of the $\bf\mu$}-metal itself, which can be machined with micrometer precision, over large scale. Mesoscopically, surface corrugations with a characteristic length scale $\xi$ of micrometers are typically present in these plates, but the associated fringe fields decay exponentially along $y$ [27, 28]. Therefore, to mitigate the effects of fringe field irregularities, 2D trapping should be positioned at a distance $y_{h}=h\gg\xi$.
@@
Figure 2: Numerical simulation of the 2D-quadruple field [29, 30] for the 2-plate assembly, (a,b), according to Sec. III.1; and 4-plate assembly, (c,d), according to Sec. IV.1. Here L=120L=120~$m$, W=40W=40~$m$, d=10d=10~$m$. See Appendix A for simulation details. The plate dimensions are not to scale. A.S.: Optional, additional shielding of source magnet variations for operating 2D-trap at large $h,h^{\prime}$ distances [31].
II.2 $n=2$ plate trap
@@
As in Fig. 1c, Fig. 2(a,b) and Fig. 3, we consider arranging $n=2$ plates symmetrically with respect to the $y=0$ plane with a gap distance of $\Delta y=2h$. With the opposite magnetization, the fields generated by the two plates cancel at ${\bf r}_{0}=(0,0,0)$ to form a magnetic zero. Near the center, the field ${\bf B}=B_{1}^{x}x{\bf e}_{x}+B_{1}^{y}y{\bf e}_{y}+B_{1}^{z}z{\bf e}_{z}$, with gradients $B_{1}^{x}:B_{1}^{y}:B_{1}^{z}=(1-\varepsilon):-1:\varepsilon$. Here $\varepsilon\sim(h/L)^{2}$ when $\varepsilon\ll 1$. The field zero is therefore a three-dimensional (3D) one. The residual $B_{1}^{z}$ gradient is associated with the fact that the field contributions from the two plates add up along $z$ for off-center locations. Nevertheless, with an operation distance $\varepsilon=h/L\ll 1$, the field distribution becomes highly elongated along $z$ to mimic a 2D field, as suggested by the dashed-line box in Fig. 2(c,d), with an approximate zero-field line near the center.
II.3 $n=4$ plate trap
@@
As in Fig. 1d and Fig. 2(c,d), we now consider arranging $n=4$ plates symmetrically in the xyx-y plane, with a gap distance of $\Delta y=2h$ and $\Delta x=2h^{\prime}$ (Fig. 2b). As in Fig. 2c, in contrast to the $n=2$ case, in the $n=4$ assembly the field variation along $z$ caused by each plate due to the finite $L$ are completely canceled at $x=y=0$. This results in a quadrupole field distribution, ${\bf B}=B_{1}y{\bf e}_{x}+B_{1}x{\bf e}_{y}$ at the center. Furthermore, with $h,h^{\prime}\ll W$, the gradient $B_{1}$ varies extremely slowly within e.g. $|z|<L/4$, as suggested by the dashed-line box in Fig. 2(c,d).
II.4 Electronic tuning of field offset
The magnetic response of $\mu$-metal is quite fast [26]. Building on earlier ferromagnetic trapping techniques [18, 19], we combine permanent magnetization with electronic tuning to generate strong and adjustable quadrupole field gradients $B_{1}$ at large distances $h$. As illustrated in Fig. 1 and demonstrated by the data in Fig. 3f, small adjustments to the currents surrounding the plates can offset the magnetization of the $\mu$-metal and shift the zero-field line. By varying the currents in selected plates of the $n$-plate assembly in Fig. 2, the position of the zero-field line can be shifted over millimeter distances in the $x$-$y$ plane within tens of microseconds, with high precision and repeatability. By aligning the local zero-field line with the ONF position, field-free operation of atomic physics or quantum optics experiments can be performed at the ONF interface. On the other hand, uniform in-plane field offsets within ten-Gauss level can be applied for e.g., canceling local vector shifts naturally exits at nanophotonic interfaces [32, 33] or simply for atomic state preparation.
Finally, while applying a uniform magnetic offset in the $x$-$y$ plane is feasible, achieving a uniform $B_{z}$ component in the $n$-plate geometry is more challenging. Even a spatially uniform external field can induce non-uniform magnetization responses in the $\mu$-metal, compromising the uniformity of the field along the $z$-axis. Nevertheless, for sufficiently large plate structures, the non-uniformity of the applied field scales with the ratio $l/L$, across the ONF interaction length $l$, and can therefore be quite negligible for most transient applications.
@@ II.4 Electronic tuning of field offset
Figure 3: (a): Schematic of the ONF-2D-MOT apparatus in this work. (b): The ONF-2D-MOT interface through the vacuum window. The ONF is visible through Rayleigh scattering of a 100 $\rm{\bf\mu}W$ heating light at 843 nm. (c): The 2D-MOT (marked with red arrow) and its three mirror images. (d): Schematic of the ONF-2D-MOT in planar geometry aided by two pairs of right-angle prisms. The dash circle represents the $x$-coil with diameter D=280D=280~$m$ which is 35 mm away from ONF. (e): Time-dependent fluorescence count through ONF and effective atom number $N_{\rm eff}$ during the MOT loading. The fit model for loading the density-limited MOT [20] (Appendix B), with $\rho_{0}=3\times 10^{10}/{\rm cm}^{3}$ final density near ONF, suggests a MOT lifetime of $\kappa^{-1}=0.4~{\rm sec}$. (f): Steady-state fluorescence count as a function of 2D-MOT position, adjusted by the $x$-coil and the plate-coil currents.
@@
For the two-plate assembly, the easiest way to form the quasi-2D-MOT is to use a right-angle prism pair that retro-reflect a single beam (Fig. 1c). The confinement along the $z$-direction is provided by two additional beams, also with two prisms in this work (Fig. 3d). This planer arrangement can be conveniently integrated with an ONF mount [13] (Fig. 3(a-c)), as suggested by Fig. 1c. By adjusting the plate-coil current to shift the field along $y$, and an $x$-coil current to shift the field along $x$, the zero-field line is precisely adjusted in the $n=2$ assembly to overlap with ONF. Clearly, the two-plate configuration only supports a quasi-2D-MOT. Depending on the $\varepsilon=h/L$ ratio, the weak gradient along the ONF direction may limit the field-free ONF-interaction length $l$.
On the other hand, as illustrated with Fig. 2(c,d), a true 2D quadruple field can be formed with the $n=4$ plate assembly. Similar to Fig. 1c, one prefers a planer geometry for the 2D-MOT construction, which can then be integrated with a standard ONF-mount [13]. Here, with the field lines rotated by $45^{\circ}$ comparing to the $n=2$ case, a 2D mirror MOT [19] is likely preferred for optimal optical accesses. The ONF can be integrated on top of the mirror along $z$. The zero-field line can be shifted freely in 2D by adjusting, e.g., the currents surrounding the plate pair on the left of Fig. 2d.